OpenStudy (anonymous):

Find integral (x/(2x^2-3)^3)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}(x \div (2x ^{2}-3)^{3}) dx\]

OpenStudy (amistre64):

whats the derivative of the bottom give us?

OpenStudy (amistre64):

prolly a bad first thought by me tho

OpenStudy (amistre64):

we can try to clean this up with a u sub

OpenStudy (anonymous):

wait, what? okay lets do a u sub

OpenStudy (amistre64):

the most likely candidate is to make u = 2x^2 - 3 so that we can just eat up that denominator

OpenStudy (amistre64):

lets take the derivative of u now so that we can find a suitable replacement for dx

OpenStudy (anonymous):

okay so with u=2x^2 - 3 the denominator is gone? okay lets find the derivative

OpenStudy (amistre64):

not gone, but cleaned up so that we can better see it

OpenStudy (anonymous):

oh okay i see what you mean

OpenStudy (amistre64):

all that stuff is hidden inside or behind the u for now

OpenStudy (amistre64):

u = 2x^2 - 3 we need to derive this for a suitabel dx replacement u' = 4x x' ... u'/4x = x' du/4x = dx is fine

OpenStudy (amistre64):

so lets start replacing stuff and simplifying

OpenStudy (anonymous):

here it is

OpenStudy (anonymous):

okay, would it be \[(-1\div4(2x ^{2}-3)^{4})+C\]

OpenStudy (amistre64):

x x du ----------- dx ; ---- ---- = \(\large \frac{1}{4u^3}du\) (2x^2-3) ^3 u^3 4x

OpenStudy (anonymous):

@harishan16 we havent learned sec and tan yet

OpenStudy (amistre64):

if you type in frac{a}{b} in the equation editor you get a prettier fraction

OpenStudy (anonymous):

dx is not scaned in that img, but still u cant undstnd!

OpenStudy (anonymous):

oh okay i was wondering how to do so

OpenStudy (amistre64):

so this cleans up to:\[\frac{1}{4}\int u^{-3}du\]

OpenStudy (amistre64):

by hiding everything we can behind us it cleans up nicely

OpenStudy (anonymous):

so I send another easy method

OpenStudy (anonymous):

my answer choices are \[\frac{-1}{2(2x ^{2}-3)^{2}} +C\] , \[\frac{-1}{4(2x ^{2}-3)^{4}}+C\] , \[\frac{-1}{16(2x ^{2}-3)^{4}}+C\]

OpenStudy (amistre64):

you have to undo the u after you integrate

OpenStudy (anonymous):

yeah @taffytwink when you sub out the u you'll end up with answer number 2

OpenStudy (amistre64):

really?

OpenStudy (anonymous):

or this answer \[\frac{-1}{8(2x ^{2}-3)^{2}}+C\]

OpenStudy (amistre64):

that one looks better to me

OpenStudy (anonymous):

my answer choices dont line up with your answer @harishan16

OpenStudy (anonymous):

sorry, u hv to divide by two

OpenStudy (amistre64):

if anything ^-3 ints up to ^-2/-2

OpenStudy (anonymous):

so which of the four answer options do you think is right

OpenStudy (amistre64):

integrate the u sub, finish it out and youll be able to pick it right out

OpenStudy (anonymous):

how do i do that

OpenStudy (amistre64):

um, the power rule for integration is the simplest thing id try ....

OpenStudy (anonymous):

alright, the answer choice doesnt go that far though so im confused

OpenStudy (anonymous):

-1/8(2x^2 -3)^2 +C

OpenStudy (anonymous):

yes harishan has the correct answer

OpenStudy (amistre64):

the real question apparently is why that would be the correct answer

OpenStudy (anonymous):

i forgot how i figured it out but it was circled so i either calculated it or guessed it right

OpenStudy (anonymous):

this is the corrected answer.

OpenStudy (amistre64):

\[\frac{1}{4}\int u^{-3}du\] is a rather basic integration and ints up with the power rule; afterwards you simply uncover the xparts behind the u

OpenStudy (anonymous):

ohhh okay, thank you so much for explaining

OpenStudy (anonymous):

heck yea @harishan16 thanks for all the help

OpenStudy (amistre64):

good luck

OpenStudy (lgbasallote):

@harishan16 you take out 1/4 not 1/2 :D the derivative of 2x^2 - 3 is 4xdx remember?

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

derivative of 2x^2 - 3 is 4xdx only if the derivative respect to x. if that respect to x^2, that is 2.

OpenStudy (lgbasallote):

ahh you took it wrt x^2 huh i see...i do wonder why the answer doesnt line up with the choices as well :/